OpenStudy (anonymous):

If the sin 90 degrees is 1 then the cos is what degrees?

OpenStudy (isaiah.feynman):

0

OpenStudy (anonymous):

\[\sin^2(90)+\cos^2(90)=1\] \[1+\cos^2(90)=1\] \[cos^2(90)=0\] \[cos(90)=0\]

OpenStudy (anonymous):

So cos (0 degrees) would be 1?

OpenStudy (isaiah.feynman):

@mustafa2014 Can you show the proof geometrically? I can't recall it

OpenStudy (isaiah.feynman):